OpenStudy (anonymous):

which is the equation of a line that passes throughthe point (3, -4) and has a slope of 2

OpenStudy (owlcoffee):

When we are given a point and a slope, we can find the equation of the line by using the following formula: \[(y- y_o )= m(x- x_o )\] where "m" is the slope we are given and "xo","yo" are the coordinates of the point we are also given.

OpenStudy (anonymous):

I understand that but still can not figure what the answer to the problem would be. Do you know what the complete answer is?

OpenStudy (owlcoffee):

Let's take the formula and set the data we have: \[(y+4)=2(x-3)\] Let's simplify it further: \[(y+4)=2x-6\] A equation of the line has several forms, but you only need to know two of them, wich are the point-slope form and the general form. The general form is when the whole equation is made equal zero: \[Ax+By+C=0\] And the point-slope form is pretty much isolating the "y" variable and the rest on the right side: \[y=mx+c\] Those are the two forms you can basically bring a line to. So let's retake the problem and say i want to bring it to it's point slope form, wich means I just want to isolate the y on the left side: \[(y+4)=2x-6\] \[y=2x -10\] And that is an answer. Why don't you try bringin it to it's general form?